# KB-DMGen: Knowledge-Based Global Guidance and Dynamic Pose Masking for Human Image Generation

### v1

@@ III-C Dynamic Masking
TABLE II: Ablation experiments on Human-Art dataset
@@ IV-A Experimental Settings
Datasets We evaluated our model on the Human-Art [34] dataset. The Human-Art dataset comprises 50,000 high-quality images from 5 real-world and 15 virtual scenarios, featuring human bounding boxes, key points and textual descriptions.
Implementation Details Similar to previous work [3, 4, 2], we fine-tune our model on SD with version 1.5. We utilize Adam [35] optimizer with a learning rate of $1\times 10^{−5}$. We also follow [3] to randomly replace text prompts as empty strings at a probability of 0.5, which aims to strengthen the control of the pose input. During inference, no text prompts are removed and a DDIM sampler [36] with 50 time steps is utilized to generate images. We train our models with 10 epochs.
For the Knowledge Base, we fine-tune on the OpenCLIP [37, 38] implementation using the pretrained CLIP model (ViT-L/14) with frozen encoders. Training is performed in two stage—codebook optimization and classifier fine-tuning—both using the same strategy over 30 epochs with AdamW (initial learning rate 0.001) and a cosine decay schedule. The training and validation splits as same as Human-Art [34].
Metrics. To evaluate pose accuracy, we use mean Average Precision (AP), Pose Cosine Similarity-based AP (CAP), and People Counting Error (PCE) [39], computed using HigherHRNet [40] to compare ground-truth poses with those extracted from generated images. For image quality assessment, we employ Fréchet Inception Distance (FID) [41] and Kernel Inception Distance (KID) [42] to measure diversity and fidelity. Text-image alignment is evaluated using the CLIP score [33].
@@ IV-B Comparison with SOTA Methods
The experimental results indicate that KB-DMGen achieved the highest AP and CLIP-score. On the Human-Art dataset, our model attained AP and CAP of 51.71% and 71.40% respectively, bringing an improvement of 2.21% in AP, compared with GRPose [15] and an improvement of 0.56% in CAP.Notably, our method significantly improves pose accuracy while ensuring the global quality (as measured by FID and KID). KID is multiplied by 100 for Human-Art.
We further compared the visualized qualitative results with other methods. As shown in Fig. 4, our method is superior in terms of image pose accuracy and the matching degree between images and text.
@@ IV-C Ablation Studies
Ablation experiments on the Human-Art dataset to validate the effectiveness of each module, as shown in Table II.
KB. When introducing the visual knowledge base, our method outperforms the recent SOTA method GRPose [15] with a 1.23 AP and 0.2 CAP improvement, while achieving better image quality with a 2.48 drop in FID and a slight 0.01 decrease in KID. This demonstrates that the KB enhances pose accuracy without compromising and even improving image quality.
DM. Building on the knowledge base, we further introduce Dynamic Masking (DM), which brings a 0.67 improvement in pose accuracy and a 0.72 reduction in FID, indicating enhanced image quality.
Hierarchical Language. To enhance the semantic parsing capability of the classifier, we decompose the language input into individual components, then query the codebook from the knowledge base (KB) for each component separately. The queried results are fused and injected into the stable diffusion (SD) model. We refer to this process as ”Decomposition and Combination (D&C)”. The results show that adding D&C based on KB and DM further improves AP and SAP, with other metrics also optimized to varying degrees. This demonstrates the effectiveness of D&C.
@@ V Conclusion
We present an image generation approach that integrates a visual knowledge base with dynamic masking. Experimental results demonstrate the effectiveness of the proposed method.

### v2

@@ III-D Loss of KB-DMGen
TABLE II: Ablation study of KB, DM, and their joint effects on Human-Art dataset.
@@
+KB w/o $\gamma_{t}$+DM w/o Sig. (KB-DMGen) 53.47 72.33 1.56 10.54 2.53 32.43
TABLE III: The KB ablation experiments on the Human-Art dataset.
[…]
Base 48.88 70.83 11.12
+KB 50.73 71.04 11.28
+KB(w/o $\mathcal{L}_{\text{entropy}}$) 50.61 71.16 11.33
+KB(w/o $\gamma_{cb}\&\beta_{cb}$) 48.34 70.57 11.70
+KB(w/o $\gamma_{t}\&\beta_{t}$) 51.12 71.70 11.85
TABLE IV: DM ablation experiments on Human-Art dataset.
[…]
Base 48.88 70.83 11.12
+DM 49.16 70.63 11.41
+DM(w/o sigmoid) 49.36 71.43 12.45
TABLE V: The number of codebook entries $K$ Ablation experiments on Human-Art dataset
[…]
256 50.09 71.06 1.56 11.34 2.47 32.48
512 50.35 70.85 1.56 11.11 2.44 32.45
1024 50.73 71.04 1.58 11.28 2.52 32.47
2048 50.40 71.08 1.52 11.49 2.53 32.47
TABLE VI: Number of trainable parameters across different stages and the number of codebook entries $K$ in KB.
Stage Trainable Parameters (M)
$K=256$ $K=512$ $K=1024$ $K=2048$
1 9.70 9.96 10.49 11.54
2 78.33 78.35 78.38 78.45
TABLE VII: Number of parameters of FiLM across different inject layers $L_{inj}$ in KB-DMGen.
[…]
Params (M) 10.8 28.2 28.2 67.3
@@ IV-A Experimental Settings
Datasets We evaluated our model on the Human-Art [39]. The Human-Art dataset comprises 50,000 high-quality images from 5 real-world and 15 virtual scenarios, featuring human bounding boxes, key points and textual descriptions. The LAION-Human dataset [1] consists of approximately 1 million image-caption pairs, filtered by high image quality and human estimation confidence scores, using a diverse range of human activities and more realistic images.
Implementation Details Similar to previous work [3, 4, 2], we fine-tune our model on SD with version 1.5. We utilize Adam [40] optimizer with a learning rate of $1\times 10^{−5}$. We also follow [3] to randomly replace text prompts as empty strings at a probability of 0.5, which aims to strengthen the control of the pose input. During inference, no text prompts are removed and a DDIM sampler [41] with 50 time steps is utilized to generate images. We train our models with 10 epochs on all datasets. The number of codebook entried $K$ is set 1024. The batch size of inference is same as Stable-Pose [2].
[…]
(a) $L_{inj}=\{4\}$
[…]
(b) $L_{inj}=\{7\}$
[…]
(c) $L_{inj}=\{10\}$
Figure 7: Visualization of the modulation parameters $\gamma$ and $\beta$ in Eq. 12 across the reverse sampling steps ($t=50\rightarrow 1$) at different injection layers ($L_{\text{inj}}=\{4,7,10\}$). (a) Shallow layers ($L_{inj}=\{4\}$) show strong and rapid variations, mainly guiding coarse structures. (b) Middle layers ($L_{inj}=\{7\}$) exhibit moderate and smoother modulation, transitioning features from coarse to fine structures. (c) Deep layers ($L_{inj}=\{10\}$) remain stable with small amplitude, mainly fine-tuning details.
[…]
Figure 8: Visualization of the DM modulation parameter $\delta$ in Eq.15 along the reverse sampling steps ($50\rightarrow 1$).
For the training of KB, we fine-tune on the OpenCLIP [42, 43] implementation using the pretrained CLIP model (ViT-L/14) with frozen encoders. Training is performed in two stage—codebook training and classifier training—both using the same strategy over 30 epochs with Adam [40] (initial learning rate 0.001), a cosine decay schedule and 256 batch size on per GPU. The training and validation splits as same as Human-Art [39]. On the Human-Art dataset, two-stage training takes about 2.5 hours with two NVIDIA 3090 GPUs.
Metrics. To evaluate pose accuracy, we use mean Average Precision (AP), Pose Cosine Similarity-based AP (CAP), and People Counting Error (PCE) [44], computed using HigherHRNet [45] to compare ground-truth poses with those extracted from generated images. For image quality assessment, we employ Fréchet Inception Distance (FID) [46] and Kernel Inception Distance (KID) [47] to measure diversity and fidelity. KID is multiplied by 100 for Human-Art.Text-image alignment is evaluated using the CLIP-score [48].
@@ IV-B Comparison with SOTA Methods
On the Human-Art dataset, our final model achieves the highest AP and CAP, reaching 53.47 and 72.33 respectively, which surpasses GRPose [15] by +3.97 AP and +1.49 CAP. Compared with the baseline StablePose [2], the improvement is even more substantial (+4.59 AP and +1.50 CAP). Meanwhile, our method maintains competitive global image quality, with FID and KID comparable to other strong baselines. This demonstrates that our design not only significantly improves pose accuracy but also preserves overall image realism. Qualitative results in Fig. 4 and Fig. 5 and further confirm that our method produces visually more faithful images with superior pose accuracy and semantic consistency.
@@ IV-C Ablation Studies
Our method is built upon Stable-Pose [2], which serves as the baseline. Ablation studies are conducted on the Human-Art dataset.
Overall Effects of KB and DM. We first evaluate the effectiveness of the proposed KB and DM modules, as well as their joint impact on performance. As shown in Table II, adding KB improves both AP and CAP ensuring image quality. It outperforms the recent SOTA method GRPose [15] with a 1.23 AP and 0.2 CAP improvement, while achieving better image quality with a 2.48 drop in FID and a slight 0.01 decrease in KID. DM alone mainly enhances AP slightly. Importantly, their combination consistently improves on AP, CAP, FID, while other indicators have only slightly decreased, demonstrating the complementarity between KB and DM. Surprisingly, when removing the temporal scaling factor $\gamma_{t}\&\beta_{t}$ from KB in Eq. 11 and the sigmoid gating from DM in Eq. 15, compared with KB+DM, this method achieves the best AP and CAP with only a slight increase in PCE, demonstrating the complementary role of KB (w/o $\gamma_{t}\&\beta_{t}$) global guidance and DM local refinement, as further confirmed by the visualizations in Sec. IV-D. What’s more, applying component-wise KB querying (+KB+DM+D&C) achieves the best overall performance compared with KB+DM. Specifically, D&C leverages the simple and regular structure of HumanArt [39] descriptions to decompose each text into three parts—type, object, and status (e.g., ”cartoon, an animal character with a sword in the woods” -¿ type = ”cartoon”; object = ”an animal character”; status = ”with a sword in the woods”.). Each component independently queries the KB, and the results are combined into the diffusion model, which enhances semantic parsing by capturing fine-grained correspondences between visual features in KB and textual components.
Effect of KB. To further investigate the KB design, we conduct several ablations as reported in Table III. Removing the entropy regularization $\mathcal{L}_{\text{entropy}}$ degrades performance in AP and FID. Eliminating the affine modulation parameters $(\gamma_{cb},\beta_{cb})$ causes a significant drop in AP and CAP, highlighting their necessity for HIG. Interestingly, removing the timestep-dependent modulation $(\gamma_{t},\beta_{t})$ yields the highest CAP and AP, but comes at the cost of worse generative fidelity (FID). We omit KID for brevity, as its variance across settings is negligible.
Effect of DM. Table IV summarizes the ablations on the DM module. While DM improves AP compared to the baseline, removing the sigmoid gating ($\delta=\text{Sigmoid}\{\textbf{MLP}(t)\}$) in Eq. 15—where the MLP has only 0.6M parameters—further increases AP and CAP but causes severe degradation of FID by about 1. This indicates that the Sig. (Sigmoid) gate is crucial for stabilizing the modulation strength and preserving global quality, even though it slightly limits precision gains.
Joint Effects of $\gamma_{t}\&\beta_{t}$ and Sigmoid. The joint analysis in Table II shows that the combination of KB without $\gamma_{t}\&\beta_{t}$ and DM without sigmoid achieves the best AP (53.47) and CAP (72.33). When used individually, each component requires additional constraints ($\gamma_{t}\&\beta_{t}$ or sigmoid) to stabilize training and prevent quality degradation, whereas their joint application exhibits complementary benefits. Specifically, KB without $\gamma_{t}\&\beta_{t}$ provides robust global guidance, while DM without sigmoid enhances local detail modulation. With global guidance in place, local refinement no longer needs the cautious restriction of sigmoid, and conversely, with flexible local control, global modulation does not require timestep-dependent $\gamma_{t}\&\beta_{t}$. Although KB and DM can also exhibit some complementarity, these constraints limit their full potential. Overall, this synergy achieves the best balance between pose precision and image quality.
Codebook Size. Finally, we analyze the impact of different codebook sizes $K$ in Table V. We observe that the performance first improves as $K$ increases, peaking at $K=1024$, and then slightly decreases with larger codebooks. This indicates that too small a codebook limits the representational capacity, while an excessively large one introduces redundancy and instability. Thus, $K=1024$ provides the best trade-off between efficiency and expressiveness.
IV-D Visualization of Injection Behaviors
Visualization of Codebook Injection. In Fig. 7, we compare the effects of KB and DM on the FiLM parameters $\gamma$ and $\beta$ at different injection layers. KB alone provides strong global trends but shows large fluctuations during sampling. Introducing DM smooths the variations of $\gamma$ and $\beta$, enabling fine-grained local refinement. Different layers exhibit distinct roles: shallow layers use larger modulation to guide coarse structure generation, while deeper layers apply smaller, more stable adjustments to refine details.
Visualization of Dynamic Mask. As shown in Fig. 8, the curve of DM w/o Sig. gradually transitions from positive to negative during the sampling process, indicating a reversal in the modulation direction of $\delta$. This leads to unstable attention on the pose regions and results in uneven modulation. In contrast, DM effectively enhances and stabilizes the magnitude of $\delta$ in DM and KB+DM. In contrast, guided by the knowledge base, KB (w/o $\gamma_{t}\&\beta_{t}$)+DM (w/o Sig.) effectively enhances $\delta$’s magnitude, significantly improving generation accuracy while ensuring overall quality and enabling the dynamic mask to stay focused on key regions even without sigmoid.
Effect on Generated Results. We further evaluate the qualitative effects of different injection strategies on the generated images (Fig. 6). In this figure, KB (w/o $\gamma_{t}\&\beta_{t}$)+DM (w/o Sig.) combination demonstrates a clear synergistic effect, producing more accurate poses and higher-quality images than either component individually. Visualizations of KB and KB (w/o $\gamma_{t}\&\beta_{t}$), and DM and DM (w/o Sig.) when used alone, as well as additional KB-DMGen generation results, are provided in the supplementary material.
IV-E Parameters Analysis
Table VI reports the number of trainable parameters when training the knowledge base (KB) alone across its two stages and for different codebook sizes $K$. Increasing $K$ slightly increases parameters in Stage 1, while Stage 2 remains largely unaffected. Table VII shows the trainable parameters of FiLM modules for different inject layers $L_{inj}$ in KB-DMGen.
@@ V Conclusion
We propose an image generation method combining a visual KB with DM, validated by experiments. Future work includes expanding the knowledge base and optimizing dynamic masking.